Write 4 equations in vertex form of parabolas that open up, down, left, and right. And tell me the vertex of each.
Answer by lwsshak3(11628)

Write 4 equations in vertex form of parabolas that open up, down, left, and right. And tell me the vertex of each.
All vertices at origin or (0,0)
parabolas open up, down, left and right respectively as follows:
x^2=4py
x^2=-4py
y^2=-4px
y^2=4px
p=distance from focus and directrix to vertex on the axis of symmetry, that is, vertex is half way between focus and directrix